OpenStudy (anonymous):

EXAM TOMORROW PLEASE HELP Find equation for the line that passes through (-5, 12) with a x-intercept of -8 I know how to find slope, but how do i find b in y=mx+b??

OpenStudy (anonymous):

what is your slope going to be?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

you put it in point slope form (thats the easiest way to do it)

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (anonymous):

y - 12 = 4(x + 5) y - 12 = 4x + 20 y = 4x + 32

OpenStudy (anonymous):

(y = mx + b)

OpenStudy (anonymous):

your y-intercept is (0, 32)

OpenStudy (anonymous):

thank you!